Operations Report – February 7th

Crew 274 Operations Report 2/7/2023

SOL: 9

Name of person filing report: Alexis Lojek

Non-nominal systems: Toilet. Robotic Observatory.

Notes on non-nominal systems: Mission Support will install a new toilet holding tank between crews. Robotic observatory is closed until further notice. Please see more detailed explanation under “summary of any observatory issues.”


Spirit rover used: Yes

Hours: 208.3

Beginning charge: 100%

Ending charge: 49%

Currently charging: Yes

Opportunity rover used: Yes

Hours: 112.4

Beginning charge: 100%

Ending charge: 72%

Currently charging: No

Curiosity rover used: Yes

Hours: 220.5

Beginning charge: 100%

Ending charge: 76%

Currently charging: No

Perseverance rover used: No

Hours: 255.2

Beginning charge: 100%

Ending charge: N/A

Currently charging: Yes

General notes on rovers: Rover Spirit was used for the morning EVA and plugged in before noon to charge as it was a beautiful, cloudless day outside. Opportunity and Curiosity were used for the afternoon EVA, but returned after noon, so were not plugged in to charge.

Summary of Hab operations: Hab is in a nominal condition.

WATER USE: Mission Support facilitated a water delivery before we were able to measure our water usage.

Water (static tank): 314 gallons remaining

Static tank pipe heater (On or off): On

Static tank heater (On or off): On

Toilet tank emptied: No

Summary of internet: Nominal

Summary of suits and radios: Suits are in nominal condition and charging. Radios were charged yesterday and are off chargers. All conditions normal.

Summary of GreenHab operations: GHO Hines watered plants from 10:27 – 11:01 pm and harvested a number of vegetables.

WATER USE: 13.74 gallons used. By the GHO’s calculations, approximately 22 gallons will be remaining at the end of the mission, after Saturday morning’s watering.

Heater: On

Supplemental light: 4 hours (10pm-2am)

Harvest: Yes. 99 grams of tomatoes, 26 grams of lettuce, 30 grams of carrots, 26 grams of spinach, and 8 grams of kale microgreens. Total: 189 grams.

Summary of ScienceDome operations: GHO Officer Tyler Hines’ experiment is continuing to be monitored; cress microgreens have been harvested. Forty Long Duration Exposure Facility seeds were relocated to the Greenhab and are sitting on a shelf in a wet paper towel in order to further support the germination process.

Dual split: On a timer from 10pm to 7am and under Mission Support supervision.

Summary of RAM operations: No operations were conducted in the RAM.

Summary of any observatory issues: One side of the dome is not opening because the straps that provide tension for the doors were wrapped around the motor. The stress on the straps is severe enough that there is no way to release that tension to try and rewrap it. The crew has been informed that the MDRS Observatory is inoperable until further notice.

Summary of health and safety issues: N/A

Questions, concerns, and requests to Mission Support: N/A

Copyright © The Mars Society. All rights reserved. | Main Site